Mary is twice as old as Helen and Sue is twice as old as Mary. In ten years their combined ages will total 58 years. How old is each at the present time?

Thanks

Answer by TimothyLamb(4379)

m = 2h
s = 2m
(m+10) + (h+10) + (s+10) = 58
---
(m+10) + (h+10) + (s+10) = 58
m + h + s = 28
2h + h + 2m = 28
2h + h + 2(2h) = 28
7h = 28
---
Answer:
h = 4
m = 8
s = 16
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php